# VL-ICL Bench: The Devil in the Details of Multimodal In-Context Learning

### v2

@@ 6 Discussion
    Subsection B.8: Self-Extend Analysis (Figure 15).
@@
    Raw results of all figures in main paper and supplementary materials from Table 7 to 67.
@@ Appendix A Implementation and Evaluation Details
VL-ICL Bench Evaluation Metrics  We use accuracy as the metric across all subsets in our benchmark. For text-to-image generation tasks, we utilize the state-of-the-art VLLM LLaVA-Next [Liu et al., 2024a] as the judge model to decide whether the generated images contain the required object or attribute.
@@ B.2 Scaling to More Shots
To examine the maximum number of shots the models can handle and whether the model can still benefit from more shots, we further increase the support set size to 16, 32, and 64 shots. We choose three models for this experiment: OpenFlamingo 9B [Awadalla et al., 2023], IDEFICS-9B-Instruct [Laurençon et al., 2023], InternLM-XComposer2 [Zhang et al., 2023], LongVA [Zhang et al., 2024b], and Mantis-Idefics2 [Jiang et al., 2024a]. These models were selected because each image they process translates to fewer tokens (Table 6) or they have a longer context window, ensuring that they do not exceed the maximum context length when evaluated with 64 shots. IDEFICS-9B-Instruct demonstrates a better scaling capability compared to other models in most of the datasets. Besides, while InternLM-XComposer2 has strong performance in a low-shot regime, the performance quickly decreases with many shots. This may be due to the mismatch between training (4096) and testing (2048) context length (Table 6) where the extrapolation of context length has been a well-known challenging task [Press et al., 2022, Liu et al., 2024b].
@@ B.3 Chain-of-Thought Prompting
To investigate whether there is any strategy that can enhance in-context learning, one straightforward method is Chain-of-Thought (CoT) prompting [Wei et al., 2022b]. CoT prompts the model to articulate its reasoning process concerning latent variables from the support set, potentially improving its learning and inference capabilities. We experiment with Qwen-VL-Chat [Bai et al., 2023] and InternLM-XComposer2 [Zhang et al., 2023] that have state-of-the-art LLMs with strong reasoning ability. Below is the specific prompt we use. {mdframed}[backgroundcolor=gray!20] [CoT Prompt]: Let’s first think step by step and analyze the relationship between the given few-shot question-answer pairs. Give reasoning rationales.
@@ B.5 Influence of Instruction Fine-tuning
We investigate how instruction-following fine-tuning affects in-context learning capabilities. We compare two model families, each with a pre-trained version and an instruction-following fine-tuned version: Qwen-VL versus Qwen-VL-Chat [Bai et al., 2023] and IDEFICS-9B versus IDEFICS-9B-Instruct [Laurençon et al., 2023]. Their performance differences are illustrated in Figure 12. Although the outcomes vary, models not fine-tuned with instructions exhibit marginally better scalability concerning the number of shots, as seen with the TextOCR dataset. Further studies are needed to understand whether instruction-following fine-tuning harm the in-context ability.
@@ Appendix D Complete Results
In this section, we show the raw results of the figures in the main text in Section D.1 to 31, and results of supplementary materials in Section D.4.
@@ D.2 Main Results
Table 23: Results of different models on Matching MiniImageNet dataset (Accuracy %).
Model 0-Shot 1-Shot 2-Shot 4-Shot 5-Shot
OpenFlamingo-9B 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00
IDEFICS-9B 50.00 $\pm$ 0.00 50.50 $\pm$ 0.35 50.83 $\pm$ 0.85 50.00 $\pm$ 0.20 49.92 $\pm$ 0.12
IDEFICS-80B 61.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00
IDEFICS2-8B 69.00 $\pm$ 0.00 50.00 $\pm$ 0.35 50.00 $\pm$ 0.54 51.42 $\pm$ 2.04 52.92 $\pm$ 1.33
Otter 48.75 $\pm$ 0.00 50.58 $\pm$ 0.31 50.92 $\pm$ 0.42 50.42 $\pm$ 0.12 49.83 $\pm$ 0.31
InternLM-XComposer2 63.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.08 $\pm$ 1.65 50.00 $\pm$ 0.00
Qwen-VL-Chat 50.50 $\pm$ 0.50 57.32 $\pm$ 1.82 55.50 $\pm$ 1.50 56.43 $\pm$ 1.17 52.82 $\pm$ 0.49
LLaVA-Next-7B 63.00 $\pm$ 0.00 50.00 $\pm$ 0.00 49.75 $\pm$ 0.00 50.00 $\pm$ 0.00 49.75 $\pm$ 0.00
Emu2-Chat 62.15 $\pm$ 3.28 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00
VILA-2.7B 59.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00
VILA-7B 77.25 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00
Mantis-Idefics2 66.00 $\pm$ 0.00 54.17 $\pm$ 0.66 55.25 $\pm$ 1.47 50.25 $\pm$ 0.20 49.75 $\pm$ 0.35
Phi3-Vision 51.75 $\pm$ 0.00 50.08 $\pm$ 0.12 50.75 $\pm$ 0.20 52.17 $\pm$ 1.25 51.25 $\pm$ 0.71
LongVA-7B 58.25 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00 50.00 $\pm$ 0.00
LLaVA-OneVision-72B 93.00 $\pm$ 0.00 87.12 $\pm$ 0.12 93.12 $\pm$ 1.38 91.62 $\pm$ 0.62 92.62 $\pm$ 0.12
InternLM-X2d5 83.75 $\pm$ 0.00 74.92 $\pm$ 0.62 84.67 $\pm$ 0.92 69.17 $\pm$ 1.01 50.17 $\pm$ 0.12
GPT4V 52.00 76.00 82.00 81.00 82.00

### v3

@@ 6 Discussion
7 Acknowledgement
Yongshuo Zong is supported by the United Kingdom Research and Innovation (grant EP/S02431X/1), UKRI Centre for Doctoral Training in Biomedical AI at the University of Edinburgh, School of Informatics. For the purpose of open access, the author has applied a creative commons attribution (CC BY) licence to any author accepted manuscript version arising.
@@ 7 Acknowledgement
    Subsection B.8: Analysis on Context Extension (Figure 15).
@@
    Subsection B.10: Impact of In-Context Demonstration Selection (Table 7 to Table 10).
    Subsection B.11: Impact of Different Model Architectures (Figure 17).
@@
    Raw results of all figures in main paper and supplementary materials from Table 11 to 70.
@@ Appendix A Implementation and Evaluation Details
VL-ICL Bench Evaluation Metrics  We use accuracy as the metric across all subsets in our benchmark. For text-to-image generation tasks, we utilize the state-of-the-art VLLM LLaVA-Next (Liu et al., 2024b) as the judge model to decide whether the generated images contain the required object or attribute.
@@
Details on Dataset Sampling and Filtering  We include additional details on how we performed dataset sampling and filtering when creating our benchmark:
  • Fast Open-Ended MiniImageNet: we take the first 200 query examples provided in the original open-ended MiniImageNet, together with their associated support examples.
  • CLEVR: we randomly sample 800 examples from the original CLEVR dataset training scenes to use as support examples, and we select 200 examples randomly from the validation split as query examples.
  • Operator Induction: we generate this dataset ourselves and we consider all single digit combinations. We use randomly selected 80 combinations of digits as the support set and 20 combinations of digits as the query set. Each combination is used with 3 different operators.
  • TextOCR: we use 800 randomly selected examples from the original TextOCR training set as support examples and 200 randomly selected examples from the validation set as query examples. Before selecting these we did filtering to ensure we use only valid texts that are not marked as rotated. To make the task manageable, we select the largest text in the image for OCR.
  • Interleaved Operator Induction: same as for the operator induction task.
  • Fast Matching MiniImageNet: this is derived from our Fast Open-Ended MiniImageNet, using the same underlying data.
  • Text-to-Image Fast MiniImageNet: we repurpose our Fast Open-Ended MiniImageNet for this task, so we use the same underlying data.
  • CoBSAT: take the first 80% of the scenarios from the original CoBSAT dataset for the support set and the remaining 20% for the query set.
  • Fast Counting: take action and colour categories from CoBSAT, using half of animals and objects for the support examples and another half for the query examples. Create composed images by repeating the image one, two, three or four times.
  • Fast Attribute Matching: take all attribute-object pairs from CoBSAT and use 30 of them as support and 20 as query sets, with no overlaps across objects between the sets.
@@ B.2 Scaling to More Shots
To examine the maximum number of shots the models can handle and whether the model can still benefit from more shots, we further increase the support set size to 16, 32, and 64 shots. We choose three models for this experiment: OpenFlamingo 9B (Awadalla et al., 2023), IDEFICS-9B-Instruct (Laurençon et al., 2023), InternLM-XComposer2 (Zhang et al., 2023a), LongVA (Zhang et al., 2024b), and Mantis-Idefics2 (Jiang et al., 2024a). These models were selected because each image they process translates to fewer tokens (Table 6) or they have a longer context window, ensuring that they do not exceed the maximum context length when evaluated with 64 shots. IDEFICS-9B-Instruct demonstrates a better scaling capability compared to other models in most of the datasets. Besides, while InternLM-XComposer2 has strong performance in a low-shot regime, the performance quickly decreases with many shots. This may be due to the mismatch between training (4096) and testing (2048) context length (Table 6) where the extrapolation of context length has been a well-known challenging task (Press et al., 2022; Liu et al., 2024c).
@@ B.3 Chain-of-Thought Prompting
To investigate whether there is any strategy that can enhance in-context learning, one straightforward method is Chain-of-Thought (CoT) prompting (Wei et al., 2022b). CoT prompts the model to articulate its reasoning process concerning latent variables from the support set, potentially improving its learning and inference capabilities. We experiment with Qwen-VL-Chat (Bai et al., 2023) and InternLM-XComposer2 (Zhang et al., 2023a) that have state-of-the-art LLMs with strong reasoning ability. Below is the specific prompt we use. {mdframed}[backgroundcolor=gray!20] [CoT Prompt]: Let’s first think step by step and analyze the relationship between the given few-shot question-answer pairs. Give reasoning rationales.
@@ B.5 Influence of Instruction Fine-tuning
We investigate how instruction-following fine-tuning affects in-context learning capabilities. We compare two model families, each with a pre-trained version and an instruction-following fine-tuned version: Qwen-VL versus Qwen-VL-Chat (Bai et al., 2023) and IDEFICS-9B versus IDEFICS-9B-Instruct (Laurençon et al., 2023). Their performance differences are illustrated in Figure 12. Although the outcomes vary, models not fine-tuned with instructions exhibit marginally better scalability concerning the number of shots, as seen with the TextOCR dataset. Further studies are needed to understand whether instruction-following fine-tuning harm the in-context ability.
@@ B.9 Average Performance over Time
B.10 Impact of In-Context Demonstration Selection
As shown in previous work (Zhang et al., 2023b; Li et al., 2024c; Chen et al., 2024; Yang et al., 2023), different in-context demonstration selection strategies can have a large impact on the final performance. To further understand this, we employ a similarity-based selection strategy (Zhang et al., 2023b; Li et al., 2024c; Chen et al., 2024; Yang et al., 2023). Specifically, for image-to-text tasks (CLEVR and TextOCR), we use CLIP image embeddings to retrieve the top-k most similar images and their corresponding text. Conversely, for text-to-image tasks (CoBSAT), we use CLIP text embeddings to retrieve the top-k most similar texts and their associated images. Additionally, we experiment with three different arrangements of the selected top-k demonstrations: (1) from the most similar one to the least similar one, followed by the query; (2) from the least similar one to the most similar one, followed by the query; and (3) a random order of the top-k demonstrations. These are compared to the original setup which used random selection.
We select several top-performing models for our experiments, with the results presented in Table 7 to 9. In each cell, the first three numbers correspond to the three arrangements of the selected top-k demonstrations (1), (2), and (3), while the final value represents the baseline performance with random selection. For both text-to-image and image-to-text tasks, similarity-based retrieval consistently enhances performance. Specifically, similarity-based selection methods have a $\sim$3improvement in peak accuracy and a $\sim$3boost in ICL efficiency compared to the baseline. These findings demonstrate that selection strategies can effectively enhance ICL tasks, highlighting a promising direction for future work aimed at improving performance across various tasks. Detailed analysis of the changes in peak accuracy and ICL efficiency is in Table 10.
Table 7: {Similarity-based Selection on CLEVR (Accuracy %).
Model 1-shot 2-shot 4-shot 8-shot
LLaVA-OneVision -7B 48.0 / 48.0 / 48.0 / 38.2 42.5 / 37.0 / 39.7 / 33.8 33.5 / 34.5 / 35.0 / 31.5 33.0 / 33.5 / 34.0 / 28.3
Phi3-Vision 42.0 / 42.0 / 42.0 / 34.5 31.5 / 29.5 / 27.5 / 25.5 29.5 / 28.5 / 32.5 / 17.7 36.5 / 36.5 / 33.7 / 18.7
Mantis-Idefics2 36.0 / 36.0 / 36.0 / 40.2 36.5 / 35.5 / 35.3 / 32.0 28.5 / 29.5 / 28.5 / 24.5 22.5 / 21.5 / 26.3 / 21.7
InternLM-X2D5 62.0 / 62.0 / 62.0 / 63.2 62.0 / 59.5 / 61.3 / 58.5 57.5 / 55.5 / 55.8 / 56.0 53.0 / 51.5 / 51.7 / 53.2
Table 8: Similarity-based Selection on TextOCR (Accuracy %).
Model 1-shot 2-shot 4-shot 8-shot
LLaVA-OneVision -7B 40.0 / 40.0 / 40.0 / 35.7 42.0 / 42.0 / 42.2 / 42.2 43.0 / 44.0 / 43.5 / 42.2 45.0 / 46.0 / 44.8 / 44.7
Phi3-Vision 32.5 / 32.5 / 32.5 / 32.2 40.0 / 39.5 / 39.0 / 38.2 42.5 / 42.0 / 42.7 / 39.7 45.0 / 43.0 / 44.7 / 41.5
Mantis-Idefics2 12.5 / 12.5 / 12.5 / 13.8 24.5 / 23.5 / 24.7 / 25.2 27.5 / 28.0 / 29.7 / 27.8 29.0 / 28.0 / 27.8 / 27.7
InternLM-X2D5 38.0 / 38.0 / 38.0 / 36.5 38.0 / 42.0 / 39.8 / 37.0 42.0 / 42.5 / 42.5 / 42.5 27.5 / 26.5 / 27.5 / 27.0
Table 9: Similarity-based Selection on CoBSAT (Accuracy %).
Model 1-shot 2-shot 4-shot 8-shot
Seed-LLama-8B 24.0 / 24.0 / 23.7 / 15.8 35.0 / 34.0 / 33.5 / 21.8 36.0 / 35.5 / 37.2 / 27.8 34.5 / 34.0 / 33.7 / 33.7
Seed-LLama-14B 27.0 / 27.0 / 27.0 / 23.0 37.0 / 34.0 / 36.5 / 33.3 43.5 / 41.5 / 42.2 / 40.8 42.0 / 43.0 / 47.8 / 43.8
Emu2-Gen 27.5 / 27.5 / 27.5 / 23.0 39.0 / 34.0 / 36.0 / 28.7 28.5 / 38.5 / 29.5 / 27.3 23.0 / 38.5 / 21.5 / 20.8
Table 10: Zero-shot accuracy, peak accuracy and ICL efficiency scores (%, $\uparrow$) of different models for the three arrangements of similarity-based selection as well as the original random selection.
Z.s. Pk. Eff.
CLEVR Count Induction
LLaVA-OneVision-7B 05.5 48.0 / 48.0 / 48.0 / 38.2 -29.6 / -29.1 / -29.9 / -24.8
Phi3-Vision 00.0 42.0 / 42.0 / 42.0 / 34.5 -31.3 / -30.6 / -31.0 / -20.4
Mantis-Idefics2 31.5 36.5 / 36.0 / 36.0 / 40.2 0-1.9 / 0-1.9 / 0-1.2 / 0-3.9
InternLM-X2d5 06.0 62.0 / 62.0 / 62.0 / 63.2 -48.6 / -47.0 / -47.5 / -47.5
TextOCR
LLaVA-OneVision-7B 39.0 45.0 / 46.0 / 44.8 / 44.7 -03.7 / -04.3 / -03.9 / -02.8
Phi3-Vision 08.0 45.0 / 43.0 / 44.7 / 41.5 -31.2 / -30.5 / -31.1 / -29.0
Mantis-Idefics2 21.0 29.0 / 28.0 / 29.7 / 27.8 -04.0 / -03.8 / -04.6 / -04.1
InternLM-X2d5 32.0 42.0 / 42.5 / 42.5 / 42.5 -04.5 / -05.2 / -05.0 / -04.2
CoBSAT
SEED-LLaMA-8B 00.5 36.0 / 35.5 / 37.2 / 33.7 -31.2 / -30.7 / -31.1 / -24.4
SEED-LLaMA-14B 05.5 43.5 / 43.0 / 47.8 / 43.8 -32.0 / -30.9 / -32.8 / -30.2
Emu2-Gen 08.7 39.0 / 38.5 / 36.0 / 28.7 -19.0 / -25.7 / -18.5 / -15.5
B.11 Impact of Different Model Architectures
We analyse the impact of different model architectures in Figure 17. More specifically we split the models into three categories: cross-attention based, LLaVA-like and GPT4V for which the architecture is unknown. To give more context, we include time in our analysis. The results indicate that the best accuracies and ICL efficiencies are obtained by LLaVA-like models (LLaVA-OneVision-72B, InternLM-X2d5, Phi3-Vision). However, more broadly there is no clear trend between cross-attention and LLaVA-like models, and both categories can obtain comparable performance. The first models were typically cross-attention based, but researchers have continued developing them also more recently.
Figure 17: Average accuracy and ICL eff. per model architecture, with information on when the model was released. The best models are LLaVA-like, but otherwise the performance is comparable.
@@ Appendix D Complete Results
In this section, we show the raw results of the figures in the main text in Section D.1 to 34, and results of supplementary materials in Section D.4.